State and prove existence and uniqueness theorem
1
Expert's answer
2020-09-13T14:58:56-0400

Existence and uniqueness theorem is the tool which makes it possible for us to conclude that there exists only one solution to a first order differential equation which satisfies a given initial condition.


Theorem. 

Let f(x,y) be a real valued function which is continuous on the rectangle 

"R=" {"(x,y); |x-x_{o}|\leq a, |y-y_{o}|\leq b" }. Assume f has a partial derivative with respect to y and that "\dfrac{\delta f}{\delta y}" is also continuous on the rectangle R. Then there exists an interval "I = [ x_{o}-h,x_{o}+h]" (with "h\leq a" ) such that the initial value problem

"\begin{cases}
 y'=f(x,y), &\\
 y(x_{o})=y_{o} & 
\end{cases}"


has a unique solution y(x) defined on the interval I.


Note that the number h may be smaller than a. In order to understand the main ideas behind this theorem, assume the conclusion is true. Then if y(x) is a solution to the initial value problem, we must have



"y(x)=y_{o}+\int^{x}_{x_{o}} f(t,y(t))dt"


It is not hard to see in fact that if a function y(x) satisfies the equation (called functional equation)


"y(x)=y_{o}+\int^{x}_{x_{o}} f(t,y(t))dt"


on an interval I, then it is solution to the initial value problem



"y'=f(x,y)"

"y(x_{o})=y_{o}"


Picard was among the first to look at the associated functional equation. The method he developed to find y is known as the method of successive approximations or Picard's iteration method. This is how it goes:


Step 1. Consider the constant function


"y_{o}(x)=y_{o}"

Step 2. Once the function "y_{n}(x)"  is known, define the function


"y_{n+1}(x)=y_{o}+\int^{x}_{x_{o}} f(t,y(t))dt"


Step 3. By induction, we generate a sequence of functions {"y_{n}(x)" }  which, under the assumptions made on f(x,y), converges to the solution y(x) of the initial value problem



"y'=f(x,y)"

"y(x_{o})=y_{o}"